You own 19% of Acme Corporation. Its earnings before taxes are $10,000,000. The corporate tax rate is 32%. The personal tax rate on dividends is 20%. Acme Corporation pays out all of its earnings as dividends. What are your after-tax earnings on the dividends? Answer to the nearest dollar.
Answer #1

Answer:

Acme Corporation's earnings before taxes = $10,000,000

Less, Taxes at 32% = 10000000 * 32% = $3,200,000

Net Income = $10,000,000 - $3,200,000 = $6,800,000

Corporation pays all of its earnings as dividends.

You own 19% of Acme Corporation.

Dividends you receive = $6,800,000 * 19% = $1,292,000

Personal tax rate on dividends = 20%

Your after tax earnings on the dividends = $1,292,000 * (1 - 20%) = $1,033,600

Your after tax earnings on the dividends = $1,033,600